# Global Geolocated Realtime Data of Interfleet Urban Transit Bus Idling

### v4

@@ Abstract
Urban transit bus idling is a contributor to ecological stress, economic inefficiency, and medically hazardous health outcomes due to emissions. The global accumulation of this frequent pattern of undesirable driving behavior is enormous. In order to measure its scale, we propose GRD-TRT-BUF-4I (“Ground Truth Buffer for Idling”) an extensible, realtime detection system that records the geolocation and idling duration of urban transit bus fleets internationally. Using live vehicle locations from General Transit Feed Specification (GTFS) Realtime, the system detects approximately 200,000 idling events per day from over 50 cities across North America, Europe, Oceania, and Asia. This realtime data was created to dynamically serve operational decision-making and fleet management to reduce the frequency and duration of idling events as they occur, as well as to capture its accumulative effects. Civil and Transportation Engineers, Urban Planners, Epidemiologists, Policymakers, and other stakeholders might find this useful for emissions modeling, traffic management, route planning, and other urban sustainability efforts at a variety of geographic and temporal scales.
1 Background & Summary
@@ 2.2 Buffering Procedure
[…]
@@ 2.3 Subsetting Procedure
$\begin{split}i&=\text{index}\\ j&=\text{attribute}\\ \end{split}$ (4)
@@
$\begin{split}h&=\text{time-horizon}\\ \end{split}$ (6)
@@
[…]
@@
[…]
@@
[…]
@@
=$\mathcal{H}=\bigl{\{}\begin{bmatrix}w_{(i;j)\,t+h}\\ \end{bmatrix}:w\in\mathcal{A}\cap\mathcal{B}\}$ (12)
@@
$\mathcal{C}$ =$=\bigl{\{}\begin{bmatrix}{}_{(i,j=1,\ldots,5)\,t+h+1}\\ \end{bmatrix}:z\in C\bigl{\}}$ (16)
@@ 2.4 Computational Procedure
Combining Buffering Procedure and Subsetting Procedure, Algorithm 1 GRD-TRT-BUF-4I (“Ground Truth Buffer for Idling") casts them into step-by-step instructions for computing realtime idling events. Details of the algorithm are provided in Supplementary Information. There are three parameters:
@@
After $r$, $h$, and $m$ are specified, an infinite loop is conducted outside the process in which Buffering Procedure and Subsetting Procedure are both initialized and computed. Additional instructions in GRD-TRT-BUF-4I from steps 10 to 22 correspond to the subset $\mathcal{H}$. Again, details of the algorithm are provided in Supplementary Information.
@@ 2.5.1 System Architecture
First, the Serve component inputs “Fleet” - the telemetry data collected from the live location of vehicles in the network of urban transit bus fleets. It outputs “$\text{Feed}_{i}$”, the GTFS Realtime feeds, and requires the open internet or “Web” as a mechanism. Note that this process is modeled for architectural context, where no explicit controls are required.
@@
The Buffer component inputs extracted data $d_{i}$ and outputs sets $A$, $B$, $C$ taken from buffer indicies $d_{0},d_{h},$$,d_{h+1}$, respectively. It uses RAM or “Memory” as a mechanism to store the contents and is controlled by $h$, as previously described in Computational Procedure. The purpose of this component is to roll the extracted data into and out of memory.
@@ 2.5.2 System Design
Modular decoupling of components into microservices allows users to directly access any phase of the ETL pipeline throughout any and all geographic regions. The design not only allows this, but strongly encourages it. To use the realtime idling event data, users access the “Subset” microservice directly, rather than through “Write” or “Store”. This is achieved through the Subset API’s and is explained later in Usage Notes.
Note that while casting the IDEF0 architecture into an ETL microservice design, the “Buffer” components are integrated within the “Subset” microservices along with parameters r, h, and m as a practical design choice. The “Write” microservices are introduced to appropriately decouple their associated function per geographic region, whereas the alternative would require a tightly coupled design with either “Subset” or “Store”. In addition, the Agency table is persisted within “Store” and is explained in the following section Data Records.
ExtractTransformLoadUserServer 1Server 2Server $n$ExtractSubsetBufferWriteServer 1Server 2Server $n$ExtractSubsetBufferWriteServer 1Server 2Server $n$ExtractExtractSubsetBufferWriteStoreAgencyEventsRegion 1Region 2Region $N$
@@
3 Data Records
There are three ways to use the data. The first is through the Realtime Responses. The second is through a historical record of realtime responses in the Relational Database. The third is through the Static Files queried from the Relational Database.
@@ 3.1 Realtime Responses
The realtime responses are JavaScript Object Notation (JSON) objects generated by the Subset microservice. Connecting to any of the Subset API’s requires access to the region’s REST API endpoint via Websocket. This is explained later in Usage Notes.
@@ 3.3 Static Files
Figure 9 exhibits the query used to create the first static file from the relational database, previously mentioned in Relational Database. The subsequent files were created using this query corresponding to their respective datetime ranges. Table 17 provides a truncated example of the files. Again, they can be downloaded in .csv format. They are the same .csv files used for evaluating data quality in the following section Technical Validation.
@@ 4.4.1 General Error
[…]
33 Exceed Min. ($<-90^{\circ}$) 0.00 % 0.00 % 0.00 %
@@
[…]
36 Exceed Min. ($<-180^{\circ}$) 0.00 % 0.00 % 0.00 %
@@ 4.4.2 Spatial Point Error
001010101020202020303030304040404050505050606060607070707080808080909090901001001001000010101010202020203030303040404040505050506060606070707070808080809090909010010010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (a): test-data-a.csvUnweightedWeighted
001010101020202020303030304040404050505050606060607070707080808080909090901001001001000010101010202020203030303040404040505050506060606070707070808080809090909010010010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (b): test-data-b.csvUnweightedWeighted
001010101020202020303030304040404050505050606060607070707080808080909090901001001001000010101010202020203030303040404040505050506060606070707070808080809090909010010010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (c): test-data-c.csvUnweightedWeighted
@@
 $s\in S:\text{K-D Tree}_{s}=KDTree(\{(\text{latitude}_{i},\text{longitude}_{i})\mid p_{i}\in s\})$ (23)
@@
[…]
@@ 4.5 Temporal Contiguity
123456789101112131415161718192021222324002,00020002{,}0002 , $0004{,}$ $0006{,}$ $0008{,}$ $00010{,}$ $0001{,}$ $0003{,}$ $0005{,}$ $0007{,}$ $0009{,}$  [^0] HourNumber of Observationstest-data-a.csvtest-data-b.csvtest-data-c.csv
@@ 4.6 Expected Duration
The global average idling duration was as expected throughout all validation periods. Exhibited in Table 36, test numbers 110 and 111 did not find any invalid values in the duration field. Test number 112 found when the minimum idling duration was adjusted to a commonly used policy definition longer than 5 minutes [8], the proportion of idling time denominated by total operational time was 38.74%. This is broadly consistent and within the 30% to 44% range from existing studies, previously mentioned in Background & Summary. Figure 12 exhibits this comparison, as well as when the minimum idling tolerance was unadjusted according to the parameterized definition $h=1$ minute.
When the minimum idling duration in test number 113 was unadjusted according to $h=1$ minute, the average (unweighted) proportion of idling increased to 56.96%. Although approximately 18% of the average proportion of idling increased as the minimum idling tolerance decreased from 5 minutes to 1 minute, it is concluded that the duration field is still reliable and valid, given its combined consistency with existing studies when adjusted to a common policy definition of idling. The material difference between tests 112 and 113 is likely not a measurement error, but rather a unique feature of this data that warrants further investigation in downstream analyses.
@@ 6 Code Availability
7 Acknowledgements
@@ 9 Competing Interests
10 Additional Information
Correspondence and requests for materials should be addressed to Nicholas Kunz.
11 Supplementary Information
@@
8:      $\mathcal{A}=\bigl{\{}[x_{(i,\,j=1,\ldots,5)\,t}]:x\in A^{n}\bigl{\}}$, =$\mathcal{B}=\bigl{\{}[y_{(i,\,j=1,\ldots,5)\,t+h}]:y\in B^{n}\bigl{\}}$

### v5

@@ Abstract
Urban transit bus idling is a contributor to ecological stress, economic inefficiency, and medically hazardous health outcomes due to emissions. The global accumulation of this frequent pattern of undesirable driving behavior is enormous. In order to measure its scale, we propose GRD-TRT-BUF-4I (“Ground Truth Buffer for Idling”) an extensible, realtime detection system that records the geolocation and idling duration of urban transit bus fleets internationally. Using live vehicle locations from General Transit Feed Specification (GTFS) Realtime, the system detects approximately 200,000 idling events per day from over 50 cities across North America, Europe, Oceania, and Asia. This realtime data was created dynamically to serve operational decision-making and fleet management to reduce the frequency and duration of idling events as they occur, as well as to capture its accumulative effects. Civil and Transportation Engineers, Urban Planners, Epidemiologists, Policymakers, and other stakeholders might find this useful for emissions modeling, traffic management, route planning, and other urban sustainability efforts at a variety of geographic and temporal scales.
1 Introduction
@@ 2.2 Buffering Procedure
Each geographic region (US East , US West , etc.) previously mentioned in Data Sources, represents an asynchronous feed, $\text{Feed}_{i}$, exhibited in Figure 4. It is constructed from its corresponding GTFS Realtime sources. As $\text{Feed}_{i}$ enters the buffer, it is stored in memory as $d_{i}$ according to indices $d_{0},d_{1},d_{2},$$d_{h+1}$ at time $t,t+1,t+2,$,$t+h+1,$ respectively.
@@ 2.3 Subsetting Procedure
[…]
@@
[…]
@@
$\begin{aligned} A^{n}&=\{x_{(i,j)\,t}:x\in A\}\\ \end{aligned}\qquad\begin{aligned} B^{n}&=\{y_{(i,j)\,t+h}:y\in B\}\end{aligned}$ (8)
@@
$\begin{aligned} A^{n}=\Biggl{\{}\begin{bmatrix}x_{(1,j)\,t}\\ x_{(2,j)\,t}\\ \vdots\\ x_{(n,j)\,t}\end{bmatrix}:x\in A\Biggl{\}}\end{aligned}\qquad\begin{aligned} B^{n}=\Biggl{\{}\begin{bmatrix}y_{(1,j)\,t+h}\\ y_{(2,j)\,t+h}\\ \vdots\\ y_{(n,j)\,t+h}\end{bmatrix}:y\in B\Biggl{\}}\end{aligned}$ (9)
@@
$\begin{aligned} \mathcal{A}&=\bigl{\{}\begin{bmatrix}x_{(i,j=1,\ldots,5)\,t}\\ \end{bmatrix}:x\in A^{n}\bigl{\}}\end{aligned}\qquad\begin{aligned} \mathcal{B}&=\bigl{\{}\begin{bmatrix}y_{(i,j=1,\ldots,5)\,t+h}\\ \end{bmatrix}:y\in B^{n}\bigl{\}}\end{aligned}$ (10)
@@
$\mathcal{H}=\bigl{\{}\begin{bmatrix}w_{(i;j)\,t+h}\\ \end{bmatrix}:w\in\mathcal{A}\cap\mathcal{B}\}$ (12)
@@
$\mathcal{C}$ $=\bigl{\{}\begin{bmatrix}{}_{(i,j=1,\ldots,5)\,t+h+1}\\ \end{bmatrix}:z\in C\bigl{\}}$ (16)
@@ 2.4 Computational Procedure
Combining Buffering Procedure and Subsetting Procedure, Algorithm 1 GRD-TRT-BUF-4I (“Ground Truth Buffer for Idling") casts them into step-by-step instructions for computing realtime idling events. Details of the algorithm are provided in Appendix. There are three parameters:
@@
After $r$, $h$, and $m$ are specified, an infinite loop is conducted outside the process in which Buffering Procedure and Subsetting Procedure are both initialized and computed. Additional instructions in GRD-TRT-BUF-4I from steps 10 to 22 correspond to the subset $\mathcal{H}$. Again, details of the algorithm are provided in Appendix.
@@ 2.5.1 System Architecture
First, the Serve component inputs “Fleet” - the telemetry data collected from the live location of vehicles in the network of urban transit bus fleets. It outputs $\text{Feed}_{i}$”, the GTFS Realtime feeds, and requires the open internet or “Web” as a mechanism. Note that this process is modeled for architectural context, where no explicit controls are required.
@@
The Buffer component inputs extracted data $d_{i}$ and outputs sets $A$, $B$, $C$ taken from buffer indicies $d_{0},d_{h},$$d_{h+1},$ respectively. It uses RAM or “Memory” as a mechanism to store the contents and is controlled by $h$, as previously described in Computational Procedure. The purpose of this component is to roll the extracted data into and out of memory.
@@ 2.5.2 System Design
Modular decoupling of components into microservices allows users to directly access any phase of the ETL pipeline throughout any and all geographic regions. The design not only allows this, but strongly encourages it. To use the realtime idling event data, users access the “Subset” microservice directly, rather than through “Write” or “Store”. This is achieved through the Subset API’s and is explained later in Data Availability.
Note that while casting the IDEF0 architecture into an ETL microservice design, the “Buffer” components are integrated within the “Subset” microservices along with parameters r, h, and m as a practical design choice. The “Write” microservices are introduced to appropriately decouple their associated function per geographic region, whereas the alternative would require a tightly coupled design with either “Subset” or “Store”. In addition, the Agency table is persisted within “Store” and is explained in the following section Records.
ExtractTransformLoadUserServer 1Server 2Server nnExtractSubsetBufferWriteServer 1Server 2Server nnExtractSubsetBufferWriteServer 1Server 2Server nnExtractExtractSubsetBufferWriteStoreAgencyEventsRegion 1Region 2Region NN
@@
3 Records
There are three ways to use the recorded data. The first is through the Realtime Responses. The second is through a historical record of realtime responses in the Relational Database. The third is through the Static Files queried from the Relational Database.
@@ 3.1 Realtime Responses
The realtime responses are JavaScript Object Notation (JSON) objects generated by the Subset microservice. Connecting to any of the Subset API’s requires access to the region’s REST API endpoint via Websocket. This is explained later in Data Availability.
@@ 3.3 Static Files
Figure 9 exhibits the query used to create the first static file from the relational database, previously mentioned in Relational Database. The subsequent files were created using this query corresponding to their respective datetime ranges. Table 17 provides a truncated example of the files. Again, they can be downloaded in .csv format. They are the same .csv files used for evaluating data quality in the following section Validation.
@@ 4.4.1 General Error
32 Exceed Max. (> $\text{ }90^{\circ}$) 0.00 % 0.00 % 0.00 %
33 Exceed Min. ($90^{\circ}$) 0.00 % 0.00 % 0.00 %
@@
35 Exceed Max. (> $\text{ }180^{\circ}$) 0.00 % 0.00 % 0.00 %
36 Exceed Min. ($180^{\circ}$) 0.00 % 0.00 % 0.00 %
@@ 4.4.2 Spatial Point Error
0101020203030404050506060707080809090100100010102020303040405050606070708080909010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (a): test-data-a.csvUnweightedWeighted
0101020203030404050506060707080809090100100010102020303040405050606070708080909010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (b): test-data-b.csvUnweightedWeighted
0101020203030404050506060707080809090100100010102020303040405050606070708080909010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (c): test-data-c.csvUnweightedWeighted
@@
 $\forall\text{ }s\in S:\text{K-D Tree}_{s}=KDTree(\{(\text{latitude}_{i},\text{longitude}_{i})\mid p_{i}\in s\})$ (23)
@@
where $C(P_{j},D_{d})$ is a counter of $P_{j}$ and $D_{d}$, 1 is an indicator function, and $dist(P_{j},s)$ is the distance between $P_{j}$ and $s$ using $\text{K-D Tree}_{s}$.
@@ 4.5 Temporal Contiguity
12345678910111213141516171819202122232402,0002{,}0004,0004{,}0006,0006{,}0008,0008{,}00010,00010{,}0001,0001{,}0003,0003{,}0005,0005{,}0007,0007{,}0009,0009{,} [^0] HourNumber of Observationstest-data-a.csvtest-data-b.csvtest-data-c.csv
@@ 4.6 Expected Duration
The global average idling duration was as expected throughout all validation periods. Exhibited in Table 36, test numbers 110 and 111 did not find any invalid values in the duration field. Test number 112 found when the minimum idling duration was adjusted to a commonly used policy definition longer than 5 minutes [8], the proportion of idling time denominated by total operational time was 38.74%. This is broadly consistent and within the 30% to 44% range from existing studies, previously mentioned in Introduction. Figure 12 exhibits this comparison, as well as when the minimum idling tolerance was unadjusted according to the parameterized definition $h=1$ minute.
When the minimum idling duration in test number 113 was unadjusted according to $h=1$ minute, the average (unweighted) proportion of idling increased to 56.96%. Although approximately 18% of the average proportion of idling increased as the minimum idling tolerance decreased from 5 minutes to 1 minute, it is concluded that the duration field is still reliable and valid, given its combined consistency with existing studies when adjusted to common policy definitions of idling. The material difference between tests 112 and 113 is likely not a measurement error, but rather a unique feature of this data that warrants further investigation in downstream analyses.
@@ 6 Code Availability
7 Conclusion
This is the first real-time detection system for urban transit bus idling measured on a global scale. Using GRD-TRT-BUF-4I, the system processes GTFS Realtime data from over 50 cities across North America, Europe, Oceania, and Asia, detecting approximately 200,000 idling events daily. A comprehensive data validation procedure was conducted using 113 individual tests across three 24-hour periods which demonstrated 93.50% spatial point accuracy and temporal contiguity exceeding 96%. The measured idling proportions were largely consistent with the existing literature when adjusted to common policy definitions of idling, confirming the reliability of the detection methodology.
Future work should focus on expanding geographic coverage, particularly in underrepresented cities and regions, as well as integrating additional vehicle telemetry data to enhance detection accuracy. The substantial difference between 1-minute and 5-minute idling thresholds observed in validation testing warrants further investigation to better understand short-duration idling patterns and their operational significance. As GTFS Realtime adoption continues to develop internationally, the system not only provides a practical implementation for detecting urban transit bus idling but also a conceptual framework for treating the issue as a coordinated global challenge rather than an isolated municipal concern.
@@ 9 Competing Interests
10 Acknowledgements
11 Appendix
@@
8:  $\mathcal{A}=\bigl{\{}[x_{(i,\,j=1,\ldots,5)\,t}]:x\in A^{n}\bigl{\}}$, $\mathcal{B}=\bigl{\{}[y_{(i,\,j=1,\ldots,5)\,t+h}]:y\in B^{n}\bigl{\}}$